A father is four times as old as his son is. Five years back he was seven times as old as his son was then, what is the present age of the father?
Answer by lwsshak3(11628)

A father is four times as old as his son is. Five years back he was seven times as old as his son was then, what is the present age of the father.
**
let x=present age of son
4x=present age of father
5 years ago:
x-5=age of son
4x-5=age of father
..
4x-5=7(x-5)
4x-5=7x-35
3x=30
x=10
4x=40
ans:
Present age of the father=40